OpenStudy (anonymous):

Solve for x to the nearest 10th ; x^3/4=129 Please help !

OpenStudy (anonymous):

what do you think?

OpenStudy (anonymous):

it's gonna be 8.0207 approx 8.0

OpenStudy (anonymous):

first make x as the subject x^3 = 129 * 4

OpenStudy (michele_laino):

hint: we have to compute the (4/3)-th power of both sides, so we can write: \[\Large \begin{gathered} {\left( {{x^{3/4}}} \right)^{4/3}} = {129^{4/3}} \hfill \\ x = \sqrt[3]{{{{129}^4}}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

Thank you guys !!! , wait @nopen what did you do after that step

OpenStudy (anonymous):

Oh ! Oops I thought it was x^3 sorry my bad

OpenStudy (michele_laino):

Thank you! :)

OpenStudy (anonymous):

@Michele_Laino is right follow what he did :)

OpenStudy (anonymous):

Thanks :)